OpenStudy (anonymous):

what is the gcf of 117 and 26 use factor tree

OpenStudy (anonymous):

117= 9*13= 3*3*13 26= 13*2 so gcf is 13

OpenStudy (anonymous):

what about 26

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

ok